. A particle moving in the xy plane undergoes a displacement (2.0i+3.0j) m as a constant force F (5.0i+ 2.0) acts on the particle. a) Calculate the magnitudes of the displacement and the force b) Calculate the work done by c) Calculate the angle between the force and displacement.


1
Expert's answer
2021-03-12T09:17:47-0500

Displacement, d="2i+3j"

Force, F= "5i+2j"


(a) Magnitude of Displacement, "|d|=\sqrt{(2)^2+(3)^2}=\sqrt{4+9}=\sqrt{13}"


Magnitude of force, "|F|=\sqrt{(5)^2+(2)^2}=\sqrt{25+4}=\sqrt{29}"


(b) Work done, "W=F.d=(2i+3j).(5i+2j)=10+6=16 joule"


(c) As we know "W=|F||d| cos\theta"


"16=\sqrt{13}.\sqrt{29} cos\theta"


"cos\theta=\dfrac{16}{\sqrt{13}.\sqrt{29}}=0.824"


"\theta=cos^{-1}(0.824)=34.5^{\circ}"